OpenStudy (anonymous):

You made two deposits to your bank account this month. One deposit was $24.61, and the second deposit was $16.49. Your balance at the end of the month is $72.31, and you made no withdrawals. Write and evaluate an expression for your balance at the beginning of the month. A $72.31 + ($24.61 – $16.49); $80.43 B $72.31 – ($24.61 – $16.49); $64.19 C $72.31 – $24.61 – $16.49; $31.21 D $72.31 + $24.61 + $16.49; $113.41

OpenStudy (abb0t):

EndingBalance - 1stDeposit - 2ndDeposit = Beginning Balance

OpenStudy (sillybutt100):

I will give you a hint all of the sign are subtract

OpenStudy (anonymous):

C) 72.31-24.61-16.49

OpenStudy (anonymous):

Thats what I got, is that correct?

OpenStudy (anonymous):

@abb0t I got C as well is that correct?

OpenStudy (abb0t):

YES!

OpenStudy (anonymous):

Thanks! Can you help me out with this problem? The temperature on a summer afternoon is 76°F. Define a variable and write an expression to find the temperature after it changes. Then evaluate your expression for a decrease of 8 degrees Fahrenheit. A. 76 + c; 76 + (–8) = 68; 68°F B. 76 – c; 76 – (–8) = 84; 84°F C. 8 – c; 8 – (–76) = 68; 68°F D. 8 + c; 8 + (–76) = 84; 84°F

OpenStudy (abb0t):

What do you think is the answer? Help us help you.

OpenStudy (anonymous):

It should be 76+c, because if the temperature decreases, the question will give you a negative value of c. So you'll get 76 + (-8) = 68.

OpenStudy (anonymous):

Thanks adit! ;3 Already gave you a medal.

OpenStudy (anonymous):

LOL its okay. :)